OpenStudy (anonymous):

A bag contains 450 blue and yellow marbles. The probability of drawing a yellow marble is 46%. How many blue marbles are in the bag?

OpenStudy (anonymous):

Solve the following for b, the number of blue marbles:\[\frac{(100-46)}{100}=\frac{\text{b}}{450}\]

OpenStudy (anonymous):

Thanks so much for giving me an equation that helped me to find the answer!!